OpenStudy (anonymous):

What is the average velocity of the particle during the first two minutes? s(t) = 16t^3 + 12t^2-144t. t>0. v(t)=48t^2+24t-144

OpenStudy (anonymous):

No on e here does velocity answers i had one question on velocity, no one answered so yaa/....

OpenStudy (turingtest):

\[v_{avg}={s(t_2)-s(t_1)\over t_2-t_1}\] so plug in\[t_2=2,t_1=0\]and you have your answer

OpenStudy (turingtest):

can you take it from here? questions?

OpenStudy (anonymous):

i did that and i kept getting the wrong answer which is why im confused :/

OpenStudy (turingtest):

let me try then and see what I get, just a sec...

OpenStudy (anonymous):

mk the answer is -56

OpenStudy (turingtest):

I got -56 (you didn't give me units)

OpenStudy (turingtest):

well, since they give us v(t) we could say\[v_{avg}=v(t_2)-v(t_1)=96-(-144)=240\]but that's pretty sneaky...

OpenStudy (anonymous):

sorry...yea thats what i kept doing. what am i doing wrogn?

OpenStudy (turingtest):

is 240 wrong as well?

OpenStudy (anonymous):

yes the only answer is -56 m/min

OpenStudy (anonymous):

ahhh i see i got the answer nowwww

OpenStudy (turingtest):

calculator error?

OpenStudy (anonymous):

could you help me with another?

OpenStudy (turingtest):

sure, repost though please

OpenStudy (turingtest):

and hook it up with a medal holmes!

OpenStudy (anonymous):

haha i was planning on it thankss!